I just cant solve this equation.... 25x^2+16=40x
Answer by stanbon(75887)

25x^2+16=40x
25x^2-40x+16 = 0
Factor:
(5x-4)^2 = 0
x = 4/5 with multiplicity two
=================================
Cheers,
Stan H.